OpenStudy (anonymous):

Find all values of x such that f(x)>0 and f(x)<0. f(x)=x^4+3x^3-4x^2

OpenStudy (anonymous):

\[f(x)=x^4+3x^3-4x^2=x^2\left(x^2+3x-4\right)=x^2(x+4)(x-1)\] This should help to see when \(f(x)\) is negative/positive.

OpenStudy (anonymous):

I factored it to x^2(x+4)(x-1). Does that make my x intercepts (-4,0),(1,0) is the correct?

OpenStudy (anonymous):

You're missing one, \((0,0)\).

OpenStudy (anonymous):

Now you just test some \(x\) values in the intervals \[(-\infty,-4)\\ (-4,0)\\ (0,1)\\ (1,\infty) \]

OpenStudy (anonymous):

Thank you, very helpful 😄

OpenStudy (anonymous):

You're welcome